Calcium hydroxide and sulfuric acid react to produce calcium sulfate and water according to the equation:

Ca(OH)2(aq) + H2SO4(aq) --> CaSO4(s) + 2 H2O(l)

What is the molarity of H2SO4 if 0.11 liters is needed to react completely with 108 grams of Ca(OH) 2?
1
Expert's answer
2019-04-18T05:04:42-0400

"m(H2SO4)= \dfrac {m(Ca(OH)2*M(H2SO4)}{M(Ca(OH)2}=\dfrac{108g*98.08g/mol}{74.09g/mol}=142.97g" "c(H2SO4)=\dfrac{142.97g}{98.08g/mol*0.11L}=13.25mol/L"